OpenStudy (faiithhh):

2.) A boat company charges a flat fee of $35.00 plus $7.25 per hour to rent a boat. Another company charges a fee of $29.00 plus $10.50 per hour to rent the same boat. please explain how i do this? i'm completely lost.

OpenStudy (faiithhh):

Using a graphing calculator, find the number of hours for which the costs are the same. Round your answer to the nearest whole hour. this is the rest of the question, sorry

OpenStudy (faiithhh):

i don't have a graphing calc btw

OpenStudy (ewzeiger77):

One second

OpenStudy (faiithhh):

okie dokie

OpenStudy (3mar):

https://www.desmos.com/calculator Take a look at that site. It is cool and interesting! I think this would help you.

OpenStudy (faiithhh):

the lines intersect at (1.8, 48.3) how would i figure out what the number of hours is when they're the same price? @3mar

OpenStudy (3mar):

Nice question!!!! Can you share your works/steps, please?

OpenStudy (faiithhh):

uhm, i went to the desmos website, and i typed the 35 + 7.25 x into the first bar, and 29 + 10.5x into the second. i scrolled up to see where the lines intersected, and found the cooridnates i listed above. I'm not sure what to do from there, though.

OpenStudy (burak):

35 + 7.25 x = 29 + 10.5 x 6 = 3.25 x x = 1.84 which means after 1.84 hours costs will be equal. 35 + 7.25 . 1.84 =48.3 (price when both are equal) if you use graph, x integer of intersection point gives the hour, y integer gives the price. Question asks for hour (x integer), so the answer is 1.8

OpenStudy (faiithhh):

ohhh! i see! so, x is after so many hours where as y is equal to the cost?

OpenStudy (burak):

oh it says round it to the nearest whole. So it's 2

OpenStudy (burak):

yes

OpenStudy (faiithhh):

thank you so much!! i get it now (:

OpenStudy (burak):

you're welcome cutie :)